OpenStudy (anonymous):

the product of 2 consecutive even integers is 2055 the product of 2 consecutive odd integers is 3076

OpenStudy (anonymous):

please show the work =]

OpenStudy (anonymous):

let x be an even integer, then the consecutive even integer is: x+2 x*(x+2) = 2055 x^2 + x = 2055

OpenStudy (anonymous):

Same thing for the other one, let y be an odd integer, then the consecutive odd integer is: y+2 y*(y+2) = 3076 y^2 + 2y = 3076

OpenStudy (anonymous):

btw the top one is: x^2 + 2x = 2055

OpenStudy (anonymous):

By the way, part one is impossible. Even * even is ALWAYS even

OpenStudy (anonymous):

so is the second part, because odd * odd is ALWAYS odd

OpenStudy (anonymous):

yes i just came up with a number b/c i dont understand consecutive even/odds and just wanted an example to be done for me Thanks!

OpenStudy (anonymous):

are u sure consecutive odd isn't x(x+1)?

OpenStudy (anonymous):

1 - odd 1+1 = 2 - even

OpenStudy (anonymous):

so both even and odd are set to x(x+2) ?

OpenStudy (anonymous):

yea 2 + 2 = 4 1 + 2 = 3 it all depends on the first number

OpenStudy (anonymous):

Oh I see, Thanks!! btw we are learning on how to solve quadratic equations!

OpenStudy (anonymous):

btw do u own a graphing calculator?

OpenStudy (anonymous):

yea TI-89 titanium

OpenStudy (anonymous):

Omg can u do me the biggest favor?

OpenStudy (anonymous):

do u know how to find 0 routes of a quadratic equation?

OpenStudy (anonymous):

what's that?

OpenStudy (anonymous):

using the graphing calculator?

OpenStudy (anonymous):

zeros? i dunno what routes are

OpenStudy (anonymous):

yeah finding the zeros with the "trace" and "zero" option

OpenStudy (anonymous):

yea sure just post the question

OpenStudy (anonymous):

Thanks so much! okay one sec

OpenStudy (anonymous):

y=x^2-6x+4 and y=3x^2+2x-2